I have the first two parts to this question but I cannot seem to get the reciprocal method amounts to come out right. The question says to round to the nearest 4 decimals. HomeLife Life Insurance Company has two service departments (actuarial and premium rating) and two production departments (ad
Answer #1

Suppose total expense of Actuarial (including share of premium) is X and total expense of premium (including Actuarial) is Y

Then,

X = 90,000 + 0.25Y........(1)

Y = 25,000 + 0.80X........(2)

Put the value of X in equation 2

Y= 25,000 + 0.80(90,000 + 0.25Y)

Y = 25,000 + 72,000 + 0.2Y

Y - 0.2Y = 97,000

0.8Y = 97,000

Y = 121,250, premium expense (including Actuarial) = 121,250

X = 90,000 + 0.25Y

X = 90,000 + 0.25(121,250)

X = 90,000 + 30,312

X = 120,312 , Actuarial expense (including premium) = 120,312

Advertising Sale
Direct cost 70,000 50,000
Actuarial 120,312 x 10% = 12,031 120,312 x 10% = 12,031
Premium 121,250 × 15% = 18,187 121,250 × 60% = 72,750
Total cost 100,218 134,781